OpenStudy (anonymous):

solve using elimination 12x+2y=90 6x+2y=60

OpenStudy (anonymous):

Multiply the second one by -1 Combine with the first...

OpenStudy (anonymous):

12x + 2y=90 -6x -2y = -60 6x = 30 x = 5 Then use x = 5 in the first equation to solve for Y

OpenStudy (anonymous):

60+2y=90